OpenStudy (anonymous):

how to use implicit differentiate for this 3=e^(xy) -2x-y?

OpenStudy (jhannybean):

\[d(3=e^{xy}-2x-y)\]\[0=e^{xy}(1\cdot y + xy') -2-y'\]\[0=ye^{xy}+xy'e^{xy}-2-y'\]\[2-ye^{xy}=y'(xe^{xy}-1)\]Now isolate the \(y'\)